9x squared = 1
the directions said simplify
: 9x squared = 1
the directions said simplify

Answer by stanbon(18754)

9x squared = 1
the directions said simplify
---------------------
9x^2-1 = 0
(3x-1)(3x+1)= 0
x = 1/3 or x = -1/3
==========================
Chers,
Stan H.